Check whether or not R is a normal subgroup of the group H = (R + Ri + Rj+ Rk,+ )where i²=j²=k² =-1,ij=-ji,jk=-kj,ki=-ik=and
( a+ bi + cj+ dk) + (a ′ + b′i + ′jc + d′ )k = (a + a′) + ( b+ b′ i) + (c + c′ j) + ( d+ d′ k) for
d,c,b,a,d',c',b',a ′∈ R.
1
Expert's answer
2020-03-17T12:25:09-0400

Clearly,"0\in\R" .let "a,b\in\R" then "a+(-b)=a-b\in\R" .

Hence ,by one step subgroup test "\R" is a subgroup of "H"

Let "r" be any element of "H."

Then, "r+\R=\R +r \space, \forall \space r \in H"

Since,"a+bi+cj+dk+x=a'+b'i+c'j+d'k+x" "\forall \space a,b,c,d" "a',b',c',d',x\in\R" .

Hence ,"\R" is a normal subgroup of "H" by the definition of normal subgroup